a Find Fand Y he as defined in Problem 1. I define a new random variable z-x-2Y. (b) Find P(X = 21Z = 0). 12 ?-
Answer #1

SOLUTION:

4(a): Following table shows the possible values of Z and P(Z-z): 0 0 0 1 P(Z-z) 1/6 1/4 1/8 1/8 1/6 1/6 0 1/6 1/4 1/8 1/8 1/6P(Z=z) 1/6 1/4 1/8 1/8 1/6 1/6 0 -2 -4 -1 -3 P(2-0, X-2) = 0 P(Z=0) = 1/6 So